OpenStudy (anonymous):

Multiply: (x + 1)(x - 2). A. x2 - 2 B. x2 - 2x - 2 C. 2x - 2 D. x2 - x - 2

OpenStudy (anonymous):

You would use FOIL, which stands for first, outer inner, last. your answer would be x^2-x-2

OpenStudy (phoenixfire):

(a+b)(c+d) = ac + ad + bc + bd then simplify. x^2 -2x + x - 2 = x^2-x-2 (D)